OpenStudy (anonymous):

on a map 15mm represent 100 km.how many mm will be in 625c?

OpenStudy (blurbendy):

@rosy do you mean 625 cm?

OpenStudy (anonymous):

i think c denote cm...actually it is also confusing me..

OpenStudy (anonymous):

we know 1cm =10mm so 625cm =625*10 mm

OpenStudy (texaschic101):

1 cm = 10 mm so 625 cm = 6250 mm 15/100 = 6250mm/x (15mm to 100km = 6250mm to x km) cross multiply 15x = (6250)(100) 15x = 625000 x = 625000/15 x = 41666 km you better get a second opinion on this

OpenStudy (anonymous):

what is the relation between cm and km? means 1 km=?cm

OpenStudy (anonymous):

you can try using a conversion factors such that c = 1 mm = 10 k = 1000  you can also try doing the same with km try to compare the 2 lengths smaller or larger try to go step by step from mm then to cm then to km then compare

OpenStudy (texaschic101):

your problem is saying 15 mm = 100 km and 1 cm = 10 mm so 625 cm = 6250 mm so, 6250mm/15 = 416.66 x 100 = 41666 does that make sense @mth3v4

OpenStudy (anonymous):

i probably skipped a 0 on the centi

OpenStudy (anonymous):

the method is done the same way :P

OpenStudy (texaschic101):

I made a proportion by putting cm on top and km on bottom

OpenStudy (anonymous):

lol i did not see i spent so much time on that 1 (i think you could tell)

OpenStudy (texaschic101):

no wait..oops...I put mm on top and km on bottom

OpenStudy (anonymous):

y does it seem that we are trying to solve the problem again ...

OpenStudy (texaschic101):

15 mm /100 km = 6250 mm / x km (6250mm = 625 cm) can it be solved that way ?

OpenStudy (anonymous):

ok lemme try to solve it

OpenStudy (texaschic101):

@Hero ..are we doing this correctly ?

OpenStudy (texaschic101):

@ganeshie8 ..are we doing this correctly ?

hero (hero):

Just convert km to cm. Then create a proportion. Not nearly as difficult as everyone is making it.

OpenStudy (texaschic101):

I converted cm to mm and used a proportion

OpenStudy (texaschic101):

does 15mm = 100 km ?

OpenStudy (texaschic101):

because in the problem, 15 mm represents 100 km

hero (hero):

It's not equal to it. It "represents" it. It's a scale conversion.

hero (hero):

15mm, in reality, is much smaller than 100km. So it can't be equal to it.

OpenStudy (texaschic101):

so you couldn't convert km to cm in this problem, could you ?

hero (hero):

Why couldn't we?

OpenStudy (texaschic101):

so your saying convert 625 cm to km ?

hero (hero):

You can do that too. Either way would work. as long as you have consistent units, you'll be able to create the proportion.

hero (hero):

@texaschic101

OpenStudy (texaschic101):

so is my proportion I posted above one of the correct ways to set it up ?

hero (hero):

The units have to be consistent. You need two mm measurements and two km measurements.

hero (hero):

In proportion to each other.

OpenStudy (texaschic101):

I believe I set it up correctly....it just look weird

hero (hero):

Well, you are correct. It does LOOK weird.

OpenStudy (texaschic101):

Because I got 6250 mm to 41666 km

hero (hero):

I looked at it again. You appear to have it.

hero (hero):

But try to express yourself more clearly next time. Use Latex or the drawing board.

OpenStudy (texaschic101):

thanks for taking the time to check it....my real face was looking like my avatar , except for the yellow

hero (hero):

Another thing is, I always set up the proportion so that units can be canceled. So I would set it up like this: \[\frac{15 \text{mm}}{6250 \text{mm}} = \frac{100 \text{km}}{x \text{km}}\]

hero (hero):

Then after cancelling units you end up with \[\frac{15}{6250} = \frac{100}{x}\]

hero (hero):

Then you can easily solve for x no problem.

OpenStudy (texaschic101):

ohhh...I see...I like that better

OpenStudy (texaschic101):

thanks.....your my hero...lol :)

hero (hero):

yw

hero (hero):

I set all of that up assuming your conversions were correct. However, think @matricked has the right approach.

hero (hero):

We're supposed to be looking for mm, not km

OpenStudy (texaschic101):

oh....I thought we were looking for km

hero (hero):

But you still converted it to km. However you already converted to mm, which is what the problem asked for.

hero (hero):

I only meant to convert to km so that you would be able to set up a proportion so solve for the unknown millimeters. I hate jumping in to a problem late.

OpenStudy (texaschic101):

I feel stupid now

hero (hero):

Don't feel stupid. Everyone makes silly mistakes every now and then.

hero (hero):

I feel silly for not realizing you had already included the converted mm value we needed to find.

OpenStudy (texaschic101):

but it says right there in the problem cm to mm.......and I had to drag km in there

hero (hero):

I try to trust that the other person if they are helping someone knows what they are doing. But I shouldn't do that.

OpenStudy (texaschic101):

why is km even in the problem then ?

hero (hero):

The thing is, I still would have converted cm to km, then set up the proportion.

OpenStudy (texaschic101):

lol...thanks again.....I think I am just tired and delirium is setting in

hero (hero):

@matricked took a short cut.

OpenStudy (anonymous):

i think you are looking for the answer in cm :P

hero (hero):

@mth3v4, what are you talking about?

OpenStudy (anonymous):

"i think"

OpenStudy (texaschic101):

I think you are looking for mm

hero (hero):

The problem clearly states what needs to be found @mth3v4

OpenStudy (anonymous):

ok my bad

OpenStudy (anonymous):

@Hero The use of mathemical knowledge should make the work easier and logical rather than the opposite

OpenStudy (texaschic101):

I feel better now...I am not the only one who is delirious...lol

OpenStudy (anonymous):

i was saved "i thought"

OpenStudy (texaschic101):

Hero saved us both

hero (hero):

@matricked, I get what you're saying, but to each his own. As long as you get to the desired result. Next time, if have something of that nature to share with me, we can discuss it via messaging.

OpenStudy (anonymous):

i meant in my extraordinarily particular common for me" case :P

OpenStudy (texaschic101):

YAWN...good night guys.....c ya tomorrow

OpenStudy (anonymous):

gn

OpenStudy (anonymous):

that was fun ^^

OpenStudy (anonymous):

look at the statement...we have to find mm not km??????

OpenStudy (anonymous):

but @hero you are letting x km.......is'nt it wrong

OpenStudy (texaschic101):

625 cm = 6250 mm